# UltraShape 1.0: High-Fidelity 3D Shape Generation via Scalable Geometric Refinement

### v1

@@ 1 Introduction
Existing watertight remeshing techniques for 3D generative models can be broadly categorized into UDF-based, visibility-check-based, and flood-fill-based approaches. UDF-based methods typically compute unsigned distance fields (UDFs) on dense voxel grids and derive pseudo-SDFs by subtracting a small offset $\epsilon$ [dora, zhang2024clay]; however, this heuristic lacks explicit sign inference, often resulting in double-layered surfaces or the erroneous removal of valid disconnected components (e.g., wheels) when filtering for the largest connected part. Alternatively, visibility-check-based methods employ ray casting to identify interior regions [zhang2024clay, step1x3d, craftsman], which effectively seal cracks and eliminate spurious internal structures but remain sensitive to occlusions and prone to high-frequency geometric noise in complex regions. Finally, flood-fill-based strategies infer signs by expanding from exterior seeds (e.g., ManifoldPlus [huang2020manifoldplus]) to generate clean, regularized surfaces. Despite their effectiveness on closed shapes, these methods rely heavily on watertight assumptions; when applied to non-watertight or self-intersecting inputs, the fill process often leaks into the interior, yielding unintended double-layered thin shells.
@@
Alongside earlier approaches such as Score Distillation Sampling [poole2022dreamfusion, wang2023prolificdreamer, chen2023fantasia3d] and Large Reconstruction Models [hong2023lrm, tang2024lgm, xu2024instantmesh], diffusion transformer (DiT [peebles2023scalable])-based methods have recently become the leading paradigm in 3D generation. They can be broadly categorized according to their underlying representations. One major line of work adopts vector set–based representations, encoding 3D shapes as a compact set of tokens, with representative methods including 3DShape2VecSet [zhang20233dshape2vecset], CLAY [zhang2024clay], Hunyuan3D 2.0 [zhao2025hunyuan3d], FlashVDM [lai2025unleashing], and TripoSG [li2025triposg]. Vector set–based methods employ a global representation. Although this allows objects to be expressed with relatively few tokens, it constrains their capacity to model fine-grained geometric details, often resulting in over-smoothed surfaces or missing local structures. In contrast, another line of work focuses on sparse voxel–based representations, as exemplified by TRELLIS [xiang2025structured], TRELLIS.2 [xiang2025native], TripoSF [he2025sparseflex], Sparc3D [li2025sparc3d], Hi3DGen [ye2025hi3dgen], and Direct3D-S2 [wu2025direct3d]. These methods allocate tokens to spatially localized regions, enabling more accurate modeling of local geometry and high-frequency details. Nevertheless, the substantially larger number of tokens introduces significant computational and memory overhead. To mitigate this issue, sparse voxel–based approaches typically rely on two-stage generation pipelines: a sparse structure is first generated, and the structured latents are then further denoised. By combining the two representations, Ultra3D [chen2025ultra3d] adopts a coarse-to-fine framework that refines the vector set-generated geometry using sparse voxel representations. More recently, LATTICE [lai2025lattice] employs a voxel set formulation, providing structured queries for vector set–based models. By introducing spatial organization into the token space, it significantly improves both scalability and geometric fidelity.
@@
Despite their advances, existing 3D data remeshing approaches and open-sourced 3D generation methods have yet to fully address the scalability challenges inherent to high-resolution 3D geometric creation. As a result, current approaches remain difficult to deploy directly in large-scale 3D production pipelines, limiting their practical applicability in industrial settings. To address the aforementioned challenges of data quality and scalable geometry generation, we present UltraShape 1.0, a 3D diffusion framework that jointly advances data curation and generative modeling. On the data side, we introduce a robust framework for watertight geometry processing and a comprehensive data filtering strategy. These components respectively resolve topological ambiguities before surface extraction and ensure high-quality data curation, ultimately boosting the quality of geometric generation. On the generation side, we adopt a two-stage coarse-to-fine strategy, where a coarse global structure is first generated and then refined via voxel-conditioned diffusion. By formulating refinement on structured voxel queries with explicit spatial encoding, the second stage decouples spatial localization from geometric detail synthesis, enabling stable training and fine-grained geometry generation at scale.
@@ 2.1 Data Curation Pipeline
Our initial data pool comes from Objaverse [objaverse], which contains about 800K 3D models across diverse categories and styles. However, after coarse inspection, we noticed unsatisfactory data quality and identified three major issues that would hinder the training: Low-Quality Geometry. A significant portion is meaningless, consists of simple primitives, and includes objects with inconsistent geometry and texture, as well as poor modeling quality. Additionally, many scanned models exhibit fragmented meshes and incomplete topologies, which are prohibitive for watertightness and result in excessively thin structures. Inconsistent Poses. 3D assets within the same categories show erratic orientations, especially for humanoid models. We found that such misalignment impedes DiT from learning coherent shape priors, thereby inducing instability during training. Intricate Internal Structures. Many hand-crafted models are manually assembled from disconnected components, leaving tiny seams and self-intersecting faces. These meshes are prone to degenerating into hollow thin shells or fragments during watertighting. Driven by these observations, we tailored a data filtering pipeline to curate high-quality 3D models for training, which consists of the following steps:
@@
    Pose Normalization. We also used VLM to remove misposed models. Inspired by [jin2025oneshot], we also trained a pose canonicalization network to detect mis-posed models and normalize them into consistent orientations.
@@ 2.2 Geometry Generation
As shown in Fig. 2, our geometry generation framework adopts a two-stage design to balance global structural coherence and fine-grained geometric detail. Specifically, we first generate a coarse representation that captures the object’s overall shape, and then refine it using voxel-based queries to synthesize detailed, high-quality geometry. At last, we also explored the potential of our method to achieve training-free 3D stylization using the proposed representation.
@@
To accomplish the refinement stage, we first analyze the underlying reasons why existing vector set–based 3D diffusion models often struggle to generate fine-grained geometric details. We identify two closely related factors that jointly contribute to this limitation:
@@
    Vector set–based methods typically rely on VAE with surface queries. As a result, the latent vectors correspond to spatial locations distributed across the entire volume, leading to an inherently large and unstructured solution space for the diffusion model during generation. The large, unstructured latent space makes the diffusion process more difficult to converge, particularly when modeling high-frequency geometric details.
@@
    The diffusion model in vector set–based approaches is required to simultaneously generate both positional information and local geometric features within the latent representation. In contrast, sparse voxel–based methods condition on fixed spatial locations and focus primarily on synthesizing local geometry. The additional burden of jointly modeling global positioning and local shape details further increases the complexity of the diffusion task in vector set–based settings.
Motivated by the above analysis and inspired by LATTICE [lai2025lattice], we design the second-stage geometry refinement module to decouple spatial localization from geometric detail synthesis explicitly. To address the first issue, we replace surface-based queries with voxel-based queries defined on a fixed-resolution voxel grid. By constraining the diffusion process to operate over a structured, discretized spatial domain, the effective solution space of the diffusion model is substantially reduced, making the training process more stable and improving convergence. To address the second issue, we adopt the coarse-to-fine refinement strategy. Specifically, the coarse geometry generated in the first stage provides explicit spatial locations for the voxel queries used in the refinement stage. These spatial coordinates are encoded into the latent representation using rotary positional embeddings (RoPE) [su2024roformer], endowing the diffusion model with explicit spatial awareness. As a result, the model is relieved from jointly modeling global positioning and local geometry, and can instead focus on synthesizing fine-grained geometric details conditioned on fixed spatial locations. Compared to directly using surface points, voxel-point queries allow the model to better refine the geometry, without strictly adhering to the Stage-1 results. This design also enhances the test time scalability of the DiT, allowing it to approach the upper limit of the VAE better.
@@
To enable voxel-based queries in the refinement stage, we adapt the shape VAE to decode geometry beyond surface-only locations. During training, we augment the original surface query points with random spatial perturbations within a bounded range, allowing the decoder to predict valid geometric information at off-surface locations. This modification equips the VAE with the ability to decode volumetric geometry using non-surface queries, which is essential for voxel-based refinement. During inference, voxel queries obtained from the first-stage coarse geometry are randomly sampled to a predefined token budget and aligned with the latent tokens used by the diffusion model. A diffusion process is then applied to iteratively denoise the latent representation. The VAE decoder decodes the denoised latent to produce an SDF field, which is queried on a regular grid. Finally, we extract the refined surface geometry using the marching cubes algorithm, yielding high-quality 3D shapes with enhanced geometric details.
As for the DiT network structure, each transformer block begins with a self-attention layer operating on the latent tokens. Spatial information is injected into each layer using RoPE. Image conditioning is incorporated through a cross-attention mechanism using DINOv2 [oquab2023dinov2] features. To prevent irrelevant background information from interfering with geometry generation, we apply an image token masking strategy that retains only informative image tokens during cross-attention. This design ensures that the diffusion model focuses on semantically relevant visual cues while maintaining robustness to background noise.
We visualize the effect of the second-stage refinement in Fig. 3. Starting from the coarse geometry generated in the first stage, the refinement process sharpens blurred edges, recovers fine-grained geometric structures, and enhances the correspondence between geometry and image conditions. The refined meshes exhibit clearer structural boundaries and richer details, demonstrating the effectiveness of the proposed coarse-to-fine generation strategy.
@@ 3.1 Implementation Details
For coarse structure generation, we directly adopt Hunyuan3D-2.1 [hunyuan3d2025hunyuan3d] as the first-stage model, as it demonstrates strong generalization across diverse object categories among existing open-source approaches. The VAE used in the refinement stage is initialized from the Hunyuan3D-2.1 VAE and fine-tuned for 55K steps with a uniform query perturbation sampled from [1/128,1/128][-1/128,1/128]. Training is conducted progressively: we first fine-tune for 40K steps with 4096 tokens, followed by 15K steps with 8192 tokens, thereby improving training stability and enabling generalization to higher token counts. The diffusion transformer (DiT) for geometry refinement is also initialized from Hunyuan3D-2.1 and fine-tuned on our dataset. We use a voxel resolution of 128 for both training and inference, and adopt a progressive multi-stage strategy that jointly increases token count and image resolution: (1) 4096 tokens at 518 resolution for 10K steps; (2) 8192 tokens at 1022 resolution for 15K steps; and (3) 10240 tokens at 1022 resolution for 60K steps. All experiments are conducted on 8 NVIDIA H20 GPUs with a batch size of 32, using 120K filtered samples from Objaverse. During inference, we use 32768 tokens and an image resolution of 1022 with token masking unless otherwise specified. We further observe that input RGBA image quality is critical: inaccurate foreground segmentation or residual background and shadow artifacts can noticeably degrade geometry quality, highlighting the importance of robust image pre-processing for image-conditioned 3D generation.
@@ 3.2 Data Watertightening
3.3 Shape Generation
We first evaluate the test-time scalability of our method, showing results in Fig. 6. As the number of latent tokens increases, reconstruction quality consistently improves, demonstrating strong potential to reconstruct high-fidelity geometries.
@@ 3.3 Shape Generation
We then evaluate the generation performance of our method and compare it with representative open-source state-of-the-art approaches. Qualitative results are shown in Figs. 7 and 8. As shown, our method produces high-quality 3D geometry with rich details and sharp structural features, while maintaining strong consistency with the input condition image. Compared to existing baselines, our results exhibit superior geometric fidelity and improved alignment between geometry and visual appearance.
We further compare our method with commercial 3D generation models, with qualitative results presented in Fig. 9. Despite being trained on only publicly available data and limited computational resources, our method achieves generation quality comparable to that of commercial systems in both geometric detail and image–geometry consistency. These results demonstrate the strong effectiveness and competitiveness of our approach.
Finally, similar to the VAE reconstruction experiments, we also investigate the test-time scaling behavior of our DiT used in geometry generation. Fig. 10 illustrates the effect of increasing the number of shape tokens and image tokens during inference. Although the model is trained with a relatively small token budget, it generalizes well to significantly larger token counts at test time, producing substantially improved geometric details and higher-quality surfaces. This result highlights the favorable test-time scalability of our generation framework.

### v2

@@ 1 Introduction
Existing watertight remeshing techniques for 3D generative models can be broadly categorized into UDF-based, visibility-check-based, and flood-fill-based approaches. UDF-based methods typically compute unsigned distance fields (UDFs) on dense voxel grids and derive pseudo-SDFs by subtracting a small offset $\epsilon$ [2, 28]; however, this heuristic lacks explicit sign inference, often resulting in double-layered surfaces or the erroneous removal of valid disconnected components (e.g., wheels) when filtering for the largest connected part. Alternatively, visibility-check-based methods employ ray casting to identify interior regions [28, 13, 12], which effectively seal cracks and eliminate spurious internal structures but remain sensitive to occlusions and prone to high-frequency geometric noise in complex regions. Finally, flood-fill-based strategies infer signs by expanding from exterior seeds (e.g., ManifoldPlus [7]) to generate clean, regularized surfaces. Despite their effectiveness on closed shapes, these methods rely heavily on watertight assumptions; when applied to non-watertight or self-intersecting inputs, the fill process often leaks into the interior, yielding unintended double-layered thin shells.
@@
Alongside earlier approaches such as Score Distillation Sampling [18, 21, 1] and Large Reconstruction Models [6, 20, 25], diffusion transformer (DiT [17])-based methods have recently become the leading paradigm in 3D generation. They can be broadly categorized according to their underlying representations. One major line of work adopts vector set–based representations, encoding 3D shapes as a compact set of tokens, with representative methods including 3DShape2VecSet [27], CLAY [28], Hunyuan3D 2.0 [29], FlashVDM [11], and TripoSG [14]. Vector set–based methods employ a global representation. Although this allows objects to be expressed with relatively few tokens, it constrains their capacity to model fine-grained geometric details, often resulting in over-smoothed surfaces or missing local structures. In contrast, another line of work focuses on sparse voxel–based representations, as exemplified by TRELLIS [24], TRELLIS.2 [23], TripoSF [5], Sparc3D [15], Hi3DGen [26], and Direct3D-S2 [22]. These methods allocate tokens to spatially localized regions, enabling more accurate modeling of local geometry and high-frequency details. Nevertheless, the substantially larger number of tokens introduces significant computational and memory overhead. To mitigate this issue, sparse voxel–based approaches typically rely on two-stage generation pipelines: a sparse structure is first generated, and the structured latents are then further denoised. By combining the two representations, Ultra3D [3] adopts a coarse-to-fine framework that refines the vector set-generated geometry using sparse voxel representations. More recently, LATTICE [10] employs a voxel set formulation, providing structured queries for vector set–based models. By introducing spatial organization into the token space, it significantly improves both scalability and geometric fidelity.
@@
Despite their advances, existing 3D data remeshing approaches and open-sourced 3D generation methods have yet to fully address the scalability challenges inherent to high-resolution 3D geometric creation. As a result, current approaches remain difficult to deploy directly in large-scale 3D production pipelines, limiting their practical applicability in industrial settings. To address the aforementioned challenges of data quality and scalable geometry generation, we present UltraShape 1.0, a 3D diffusion framework that jointly advances data curation and generative modeling. On the data side, we introduce a robust framework for watertight geometry processing and a comprehensive data filtering strategy. These components respectively resolve topological ambiguities before surface extraction and ensure high-quality data curation, ultimately boosting the quality of geometric generation. On the generation side, we adopt a two-stage coarse-to-fine strategy following LATTICE [10], where a coarse global structure is first generated and then refined via voxel-conditioned diffusion. By formulating refinement on structured voxel queries with explicit spatial encoding, the second stage decouples spatial localization from geometric detail synthesis, enabling stable training and fine-grained geometry generation at scale.
@@ 2.1 Data Curation Pipeline
Our initial data pool comes from Objaverse [4], which contains about 800K 3D models across diverse categories and styles. However, after coarse inspection, we noticed unsatisfactory data quality and identified three major issues that would hinder the training: Low-Quality Geometry. A significant portion is meaningless, consists of simple primitives, and includes objects with inconsistent geometry and texture, as well as poor modeling quality. Additionally, many scanned models exhibit fragmented meshes and incomplete topologies, which are prohibitive for watertightness and result in excessively thin structures. Inconsistent Poses. 3D assets within the same categories show erratic orientations, especially for humanoid models. We found that such misalignment impedes DiT from learning coherent shape priors, thereby inducing instability during training. Intricate Internal Structures. Many hand-crafted models are manually assembled from disconnected components, leaving tiny seams and self-intersecting faces. These meshes are prone to degenerating into hollow thin shells or fragments during watertighting. Driven by these observations, we tailored a data filtering pipeline to curate high-quality 3D models for training, which consists of the following steps:
@@
    Pose Normalization. We also used VLM to remove misposed models. Inspired by [9], we also trained a pose canonicalization network to detect mis-posed models and normalize them into consistent orientations.
@@ 2.2 Geometry Generation
Following LATTICE [10], our geometry generation framework adopts a two-stage coarse-to-fine design to balance global structural coherence and fine-grained geometric detail. We refer readers to [10] for a comprehensive discussion of the overall formulation, and provide a brief description here for completeness. Specifically, we first generate a coarse representation that captures the overall shape of the object, and then refine it using voxel-based queries to synthesize detailed, high-quality geometry. In addition, we explore the potential of the proposed representation for training-free 3D stylization.
@@
LATTICE [10] has analyzed why existing vector set–based 3D diffusion models often struggle to generate fine-grained geometric details. In particular, two closely related factors have been identified as key contributors to this limitation:
@@
    Vector set–based methods typically rely on VAEs with surface queries. As a result, latent vectors correspond to spatial locations distributed across the entire volume, leading to an inherently large and unstructured solution space for the diffusion model during generation. Such a large, unstructured latent space makes the diffusion process more difficult to converge, especially when modeling high-frequency geometric details.
@@
    In vector set–based approaches, the diffusion model is required to simultaneously generate both positional information and local geometric features within the latent representation. In contrast, sparse voxel–based methods condition on fixed spatial locations and primarily focus on synthesizing local geometry. This coupling of global positioning and local shape synthesis further increases the complexity of the diffusion task in vector set–based settings.
With the analysis described above, the refinement stage is formulated to explicitly decouple spatial localization from geometric detail synthesis. This is achieved by performing diffusion-based refinement on voxel queries defined over a fixed-resolution grid, which constrains generation to a structured and discretized spatial domain. As a result, the diffusion process operates in a substantially reduced and more stable solution space. In this coarse-to-fine formulation, coarse geometry provides explicit spatial anchors for refinement. Voxel queries derived from the coarse shape define fixed spatial locations, whose coordinates are encoded using rotary positional embeddings (RoPE) [19]. With spatial localization explicitly specified, the diffusion model focuses on synthesizing local geometric details rather than jointly modeling global positioning and shape, leading to improved convergence and finer geometric refinement. Compared to surface-based queries, voxel-point queries allow further geometric refinement without strictly adhering to the coarse input.
@@
To support voxel-based refinement, the shape VAE is extended to decode geometry at off-surface locations. During training, surface queries are augmented with bounded spatial perturbations, enabling the decoder to predict valid volumetric geometry. At inference time, voxel queries sampled from the coarse geometry are aligned with latent tokens and refined through a diffusion process. The denoised latent representation is decoded into an SDF field on a regular grid, from which the final surface is extracted using marching cubes. The refinement stage employs a DiT architecture with self-attention over latent tokens, and the results are shown in Fig. 3. Spatial information is injected via RoPE at each layer, while image conditioning is incorporated through cross-attention using DINOv2 [16] features. An image token masking strategy is applied to suppress irrelevant background information, ensuring robust and semantically aligned geometry refinement.
@@ 3.1 Implementation Details
For coarse structure generation, we directly adopt Hunyuan3D-2.1 [8] as the first-stage model, as it demonstrates strong generalization across diverse object categories among existing open-source approaches. The VAE used in the refinement stage is initialized from the Hunyuan3D-2.1 VAE and fine-tuned for 55K steps with a uniform query perturbation sampled from [1/128,1/128][-1/128,1/128]. Training is conducted progressively: we first fine-tune for 40K steps with 4096 tokens, followed by 15K steps with 8192 tokens, thereby improving training stability and enabling generalization to higher token counts. The diffusion transformer (DiT) for geometry refinement is also initialized from Hunyuan3D-2.1 and fine-tuned on our dataset. We use a voxel resolution of 128 for both training and inference, and adopt a progressive multi-stage strategy that jointly increases token count and image resolution: (1) 4096 tokens at 518 resolution for 10K steps; (2) 8192 tokens at 1022 resolution for 15K steps; and (3) 10240 tokens at 1022 resolution for 60K steps. All experiments are conducted on 8 NVIDIA H20 GPUs with a batch size of 32, using 120K filtered samples from Objaverse. During inference, we use 32768 tokens and an image resolution of 1022 with token masking unless otherwise specified. We further observe that input RGBA image quality is critical: inaccurate foreground segmentation or residual background and shadow artifacts can noticeably degrade geometry quality, highlighting the importance of robust image pre-processing for image-conditioned 3D generation.
@@ 3.2 Data Watertightening
3.3 Shape Generation
We first evaluate the test-time scalability of our model, showing results in Fig. 9. As the number of latent tokens increases, reconstruction quality consistently improves, demonstrating strong potential to reconstruct high-fidelity geometries.
We then evaluate the generation performance of our model and compare it with representative open-source state-of-the-art approaches. Qualitative results are shown in Figs. 6 and 7. As shown, our model produces high-quality 3D geometry with rich details and sharp structural features, while maintaining strong consistency with the input condition image. Compared to existing baselines, our results exhibit superior geometric fidelity and improved alignment between geometry and visual appearance.
We further compare our model with commercial 3D generation models, with qualitative results presented in Fig. 8. Despite being trained on only publicly available data and limited computational resources, our model achieves generation quality comparable to that of commercial systems in both geometric detail and image–geometry consistency. These results demonstrate the strong effectiveness and competitiveness of our approach.
Finally, similar to the VAE reconstruction experiments, we also investigate the test-time scaling behavior of the DiT used in geometry generation. Fig. 10 illustrates the effect of increasing the number of shape tokens and image tokens during inference. Although the model is trained with a relatively small token budget, it generalizes well to significantly larger token counts at test time, producing substantially improved geometric details and higher-quality surfaces. This result highlights the favorable test-time scalability of the generation framework.